OpenStudy (anonymous):

pt. on graph of f(x)=x^2-1 closest to (-4,0)? i'm getting non-real #'s when solving d/dx for 0

OpenStudy (anonymous):

let me know if you are still here

OpenStudy (anonymous):

yes, on 2nd effort with a little help I got (√5/2,3/2)?

OpenStudy (anonymous):

best method to use is lagrange method

OpenStudy (anonymous):

i used distance formula plugging in pts. and f(x)=y=(x^2-1), then derive and set to zero?

OpenStudy (anonymous):

F= (x+4)^2 + y^2 \[\nabla F=<2(x+4),2y>\] \[\nabla F=<-2x,1>\]

OpenStudy (anonymous):

second one should have been \[\nabla G=<-2x,1>\]

OpenStudy (anonymous):

\[2(x+4)=-\lambda 2x\] \[2y=\lambda\] y-x^2=1

OpenStudy (anonymous):

gotta solve for x and y

OpenStudy (anonymous):

do my distance formula/derive/minimize results look correct?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

mahalo

OpenStudy (anonymous):

I got (0,-1)

OpenStudy (anonymous):

distance of 4.12

OpenStudy (anonymous):

(0,-1) is the vertex, no? concave up parabola?

OpenStudy (anonymous):

you numbers \[\sqrt{\left(\frac{3}{2}\right)^2+\left(\frac{\sqrt{5}}{2}+4\right)^2}\] =5.33

OpenStudy (anonymous):

y=x^2 -1 (0,-1) is on the curve y=0^2 -1 y=-1 it would be concave up

OpenStudy (anonymous):